Given that the set S= ((1,0,0,0), (0,0,1.0), (5,1,11,0), (-4,0,-6,1)) is a basis of R*, and T= {(1,0,1,0), (0,2,0,3)} is linearly independent. Extend T to a basis of R.

1
Expert's answer
2022-05-05T06:30:08-0400

At first, we compute the determinant: "A=\left(\begin{array}{llll}1&0&0&0\\0&0&1&0\\5&1&11&0\\-4&0&-6&1\end{array}\right)". We receive: "\det(A)=-1". Thus, the set "S" indeed contains the basis (the vectors of "S" are linearly independent and "5" vectors in "\mathbb{R}^4" are always linearly dependent). Another basis, which can be received as an extension of "T", can be found in different ways. For example, we can use Gram-Schmidt orthogonalization. Another way is to consider the matrix: "B=\left(\begin{array}{llll}1&0&1&0\\0&2&0&3\end{array}\right)". As we can see, the block "\left(\begin{array}{llll}1&0\\0&2\end{array}\right)" is non-degenerate. It means, that we can easily extend the matrix: "\tilde{B}=\left(\begin{array}{llll}1&0&1&0\\0&2&0&3\\0&0&1&0\\0&0&0&1\end{array}\right)". The block "\left(\begin{array}{ll}1&0\\0&1\end{array}\right)" is also non-degenerate. "\det(\tilde{B})=2". Therefore, the respective vectors of matrix "\tilde{B}" are linearly independent. Thus, they are the basis.

Answer: "(1,0,1,0)", "(0,2,0,3)", "(0,0,1,0)" and "(0,0,0,1)" is the basis. It is received as the extension of "T".